# Tuning LLMs with Contrastive Alignment Instructions for Machine Translation in Unseen, Low-resource Languages

### v1

@@ Appendix G Per Language Result Details of BLOOMZ+24 and BLOOMZ+3
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET |
| ar-de | 5.1 | 20.8 | 55.25 | en-ar | 8.4 | 26.0 | 70.45 |
| ar-fr | 20.3 | 42.5 | 71.78 | en-de | 21.1 | 36.7 | 67.15 |
| ar-nl | 6.4 | 21.6 | 57.48 | en-fr | 22.9 | 44.4 | 74.67 |
| ar-ru | 5.2 | 21.5 | 55.51 | en-nl | 16.1 | 36.8 | 65.26 |
| ar-zh | 16.0 | 14.1 | 69.55 | en-ru | 15.2 | 31.5 | 63.30 |
| de-ar | 2.4 | 16.3 | 64.01 | en-zh | 16.1 | 15.0 | 71.93 |
| de-fr | 13.5 | 34.3 | 66.25 | avg. | 16.63 | 31.73 | 68.79 |
| de-nl | 9.7 | 28.0 | 55.00 | ||||
| de-ru | 5.3 | 19.6 | 55.61 | ||||
| de-zh | 7.2 | 7.3 | 60.64 | ||||
| fr-ar | 10.0 | 28.2 | 69.86 | BLEU | chrF++ | COMET | |
| fr-de | 9.2 | 27.8 | 54.03 | ar-en | 27.1 | 47.0 | 76.54 |
| fr-nl | 10.8 | 31.0 | 58.50 | de-en | 27.8 | 44.4 | 77.57 |
| fr-ru | 8.6 | 26.7 | 57.07 | fr-en | 27.1 | 48.7 | 78.82 |
| fr-zh | 15.9 | 15.8 | 70.78 | nl-en | 22.6 | 42.2 | 74.25 |
| nl-ar | 2.2 | 15.4 | 63.47 | ru-en | 25.6 | 44.2 | 74.46 |
| nl-de | 10.2 | 28.5 | 53.65 | zh-en | 23.5 | 45.7 | 77.04 |
| nl-fr | 14.4 | 34.4 | 66.55 | avg. | 25.62 | 45.37 | 76.45 |
| nl-ru | 5.3 | 19.3 | 55.53 | ||||
| nl-zh | 5.5 | 6.2 | 58.77 | ||||
| ru-ar | 6.5 | 16.0 | 62.69 | ||||
| ru-de | 6.1 | 24.3 | 52.89 | ||||
| ru-fr | 18.2 | 39.0 | 69.95 | ||||
| ru-nl | 6.3 | 22.5 | 54.36 | ||||
| ru-zh | 7.6 | 13.3 | 61.94 | ||||
| zh-ar | 8.7 | 26.5 | 70.88 | ||||
| zh-de | 3.0 | 19.5 | 50.82 | ||||
| zh-fr | 17.7 | 39.7 | 73.56 | ||||
| zh-nl | 4.4 | 19.3 | 54.20 | ||||
| zh-ru | 4.1 | 19.5 | 50.47 | ||||
| avg. | 8.86 | 23.30 | 60.70 | ||||
| seenseen | 14.77 | 27.80 | 71.07 | enseen | 15.80 | 28.47 | 72.35 |
| seenunseen | 6.31 | 23.08 | 54.81 | enunseen | 17.47 | 35.00 | 65.24 |
| unseenseen | 8.61 | 20.24 | 63.81 | seenen | 25.90 | 47.13 | 77.47 |
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET |
| ar-de | 6.9 | 24.7 | 58.10 | en-ar | 14.6 | 35.6 | 76.70 |
| ar-fr | 26.2 | 48.2 | 74.96 | en-de | 20.4 | 36.0 | 65.96 |
| ar-nl | 8.8 | 24.7 | 59.53 | en-fr | 27.9 | 50.0 | 77.65 |
| ar-ru | 6.5 | 22.7 | 55.33 | en-nl | 14.8 | 34.8 | 63.11 |
| ar-zh | 28.6 | 22.3 | 77.64 | en-ru | 13.3 | 29.0 | 61.43 |
| de-ar | 3.3 | 19.8 | 68.27 | en-zh | 36.1 | 27.7 | 80.31 |
| de-fr | 15.2 | 35.8 | 67.05 | avg. | 21.18 | 35.52 | 70.86 |
| de-nl | 8.2 | 26.0 | 53.35 | ||||
| de-ru | 4.4 | 17.9 | 54.79 | ||||
| de-zh | 12.0 | 9.9 | 65.20 | ||||
| fr-ar | 14.2 | 35.2 | 74.84 | BLEU | chrF++ | COMET | |
| fr-de | 8.9 | 28.4 | 53.81 | ar-en | 33.7 | 53.5 | 79.81 |
| fr-nl | 10.1 | 29.9 | 56.92 | de-en | 27.1 | 43.9 | 77.04 |
| fr-ru | 8.1 | 26.0 | 55.96 | fr-en | 29.6 | 51.0 | 79.60 |
| fr-zh | 30.2 | 25.6 | 79.43 | nl-en | 22.0 | 41.4 | 73.54 |
| nl-ar | 3.1 | 18.2 | 67.72 | ru-en | 25.1 | 43.9 | 74.05 |
| nl-de | 10.4 | 27.7 | 52.67 | zh-en | 32.6 | 54.3 | 79.75 |
| nl-fr | 16.9 | 37.3 | 68.46 | avg. | 28.35 | 48.00 | 77.30 |
| nl-ru | 4.8 | 17.8 | 54.71 | ||||
| nl-zh | 8.1 | 7.0 | 63.96 | ||||
| ru-ar | 11.9 | 31.5 | 72.45 | ||||
| ru-de | 6.1 | 23.7 | 52.74 | ||||
| ru-fr | 21.2 | 42.5 | 71.71 | ||||
| ru-nl | 6.8 | 22.6 | 53.91 | ||||
| ru-zh | 21.3 | 20.7 | 74.63 | ||||
| zh-ar | 13.1 | 34.1 | 74.92 | ||||
| zh-de | 4.1 | 22.3 | 52.13 | ||||
| zh-fr | 23.8 | 46.5 | 76.54 | ||||
| zh-nl | 4.8 | 19.9 | 54.26 | ||||
| zh-ru | 5.7 | 21.9 | 50.60 | ||||
| avg. | 11.79 | 26.36 | 63.22 | ||||
| seenseen | 22.68 | 35.32 | 76.39 | enseen | 26.20 | 37.77 | 78.22 |
| seenunseen | 7.10 | 24.50 | 55.18 | enunseen | 16.17 | 33.27 | 63.50 |
| unseenseen | 12.56 | 24.74 | 68.83 | seenen | 31.97 | 52.93 | 79.72 |
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET |
| ar-de | 6.7 | 24.2 | 57.45 | en-ar | 15.1 | 35.8 | 76.76 |
| ar-fr | 27.5 | 49.2 | 75.21 | en-de | 20.6 | 35.9 | 65.88 |
| ar-nl | 8.7 | 24.8 | 59.14 | en-fr | 27.5 | 49.4 | 77.46 |
| ar-ru | 6.7 | 21.6 | 55.04 | en-nl | 15.0 | 35.6 | 63.70 |
| ar-zh | 30.1 | 24.4 | 78.54 | en-ru | 13.5 | 29.5 | 61.62 |
| de-ar | 3.5 | 19.7 | 68.39 | en-zh | 36.3 | 27.7 | 80.52 |
| de-fr | 15.4 | 35.8 | 67.81 | avg. | 21.33 | 35.65 | 70.99 |
| de-nl | 9.6 | 27.3 | 53.74 | ||||
| de-ru | 4.7 | 17.9 | 54.23 | ||||
| de-zh | 12.0 | 9.9 | 65.40 | ||||
| fr-ar | 14.9 | 36.3 | 74.98 | BLEU | chrF++ | COMET | |
| fr-de | 9.2 | 28.3 | 52.96 | ar-en | 33.9 | 53.7 | 79.74 |
| fr-nl | 11.3 | 31.1 | 57.62 | de-en | 27.1 | 43.6 | 77.13 |
| fr-ru | 8.8 | 26.2 | 56.31 | fr-en | 29.7 | 51.0 | 80.03 |
| fr-zh | 31.1 | 26.9 | 79.93 | nl-en | 22.6 | 42.3 | 73.94 |
| nl-ar | 3.3 | 18.5 | 68.02 | ru-en | 25.8 | 44.5 | 74.07 |
| nl-de | 9.4 | 26.5 | 52.33 | zh-en | 32.5 | 54.5 | 80.01 |
| nl-fr | 17.2 | 37.3 | 68.38 | avg. | 28.60 | 48.27 | 77.49 |
| nl-ru | 4.4 | 17.1 | 53.63 | ||||
| nl-zh | 8.3 | 7.0 | 64.08 | ||||
| ru-ar | 12.4 | 32.1 | 72.40 | ||||
| ru-de | 5.7 | 22.9 | 51.90 | ||||
| ru-fr | 21.5 | 42.7 | 72.08 | ||||
| ru-nl | 6.3 | 22.1 | 53.32 | ||||
| ru-zh | 22.7 | 24.6 | 75.36 | ||||
| zh-ar | 13.9 | 35.4 | 75.68 | ||||
| zh-de | 3.6 | 21.3 | 51.32 | ||||
| zh-fr | 24.5 | 47.0 | 76.98 | ||||
| zh-nl | 4.9 | 20.3 | 54.30 | ||||
| zh-ru | 5.5 | 21.1 | 50.49 | ||||
| avg. | 12.13 | 26.65 | 63.23 | ||||
| seenseen | 23.67 | 36.53 | 76.89 | enseen | 26.30 | 37.63 | 78.25 |
| seenunseen | 7.27 | 24.32 | 54.96 | enunseen | 16.37 | 33.67 | 63.73 |
| unseenseen | 12.92 | 25.29 | 69.10 | seenen | 32.03 | 53.07 | 79.93 |

### v2

@@ Appendix I Per Language Result Details of BLOOMZ+24 and BLOOMZ+3
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 5.1 | 20.8 | 55.25 | en-ar | 8.4 | 26.0 | 70.45 | |
| ar-fr | 20.3 | 42.5 | 71.78 | en-de | 21.1 | 36.7 | 67.15 | |
| ar-nl | 6.4 | 21.6 | 57.48 | en-fr | 22.9 | 44.4 | 74.67 | |
| ar-ru | 5.2 | 21.5 | 55.51 | en-nl | 16.1 | 36.8 | 65.26 | |
| ar-zh | 16.0 | 14.1 | 69.55 | en-ru | 15.2 | 31.5 | 63.30 | |
| de-ar | 2.4 | 16.3 | 64.01 | en-zh | 16.1 | 15.0 | 71.93 | |
| de-fr | 13.5 | 34.3 | 66.25 | avg. | 16.63 | 31.73 | 68.79 | |
| de-nl | 9.7 | 28.0 | 55.00 | |||||
| de-ru | 5.3 | 19.6 | 55.61 | |||||
| de-zh | 7.2 | 7.3 | 60.64 | |||||
| fr-ar | 10.0 | 28.2 | 69.86 | BLEU | chrF++ | COMET | ||
| fr-de | 9.2 | 27.8 | 54.03 | ar-en | 27.1 | 47.0 | 76.54 | |
| fr-nl | 10.8 | 31.0 | 58.50 | de-en | 27.8 | 44.4 | 77.57 | |
| fr-ru | 8.6 | 26.7 | 57.07 | fr-en | 27.1 | 48.7 | 78.82 | |
| fr-zh | 15.9 | 15.8 | 70.78 | nl-en | 22.6 | 42.2 | 74.25 | |
| nl-ar | 2.2 | 15.4 | 63.47 | ru-en | 25.6 | 44.2 | 74.46 | |
| nl-de | 10.2 | 28.5 | 53.65 | zh-en | 23.5 | 45.7 | 77.04 | |
| nl-fr | 14.4 | 34.4 | 66.55 | avg. | 25.62 | 45.37 | 76.45 | |
| nl-ru | 5.3 | 19.3 | 55.53 | |||||
| nl-zh | 5.5 | 6.2 | 58.77 | |||||
| ru-ar | 6.5 | 16.0 | 62.69 | |||||
| ru-de | 6.1 | 24.3 | 52.89 | |||||
| ru-fr | 18.2 | 39.0 | 69.95 | |||||
| ru-nl | 6.3 | 22.5 | 54.36 | |||||
| ru-zh | 7.6 | 13.3 | 61.94 | |||||
| zh-ar | 8.7 | 26.5 | 70.88 | |||||
| zh-de | 3.0 | 19.5 | 50.82 | |||||
| zh-fr | 17.7 | 39.7 | 73.56 | |||||
| zh-nl | 4.4 | 19.3 | 54.20 | |||||
| zh-ru | 4.1 | 19.5 | 50.47 | |||||
| avg. | 8.86 | 23.30 | 60.70 | |||||
| seenseen | 14.77 | 27.80 | 71.07 | enseen | 15.80 | 28.47 | 72.35 | |
| seenunseen | 6.31 | 23.08 | 54.81 | enunseen | 17.47 | 35.00 | 65.24 | |
| unseenseen | 8.61 | 20.24 | 63.81 | seenen | 25.90 | 47.13 | 77.47 | |
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 6.9 | 24.7 | 58.10 | en-ar | 14.6 | 35.6 | 76.70 | |
| ar-fr | 26.2 | 48.2 | 74.96 | en-de | 20.4 | 36.0 | 65.96 | |
| ar-nl | 8.8 | 24.7 | 59.53 | en-fr | 27.9 | 50.0 | 77.65 | |
| ar-ru | 6.5 | 22.7 | 55.33 | en-nl | 14.8 | 34.8 | 63.11 | |
| ar-zh | 28.6 | 22.3 | 77.64 | en-ru | 13.3 | 29.0 | 61.43 | |
| de-ar | 3.3 | 19.8 | 68.27 | en-zh | 36.1 | 27.7 | 80.31 | |
| de-fr | 15.2 | 35.8 | 67.05 | avg. | 21.18 | 35.52 | 70.86 | |
| de-nl | 8.2 | 26.0 | 53.35 | |||||
| de-ru | 4.4 | 17.9 | 54.79 | |||||
| de-zh | 12.0 | 9.9 | 65.20 | |||||
| fr-ar | 14.2 | 35.2 | 74.84 | BLEU | chrF++ | COMET | ||
| fr-de | 8.9 | 28.4 | 53.81 | ar-en | 33.7 | 53.5 | 79.81 | |
| fr-nl | 10.1 | 29.9 | 56.92 | de-en | 27.1 | 43.9 | 77.04 | |
| fr-ru | 8.1 | 26.0 | 55.96 | fr-en | 29.6 | 51.0 | 79.60 | |
| fr-zh | 30.2 | 25.6 | 79.43 | nl-en | 22.0 | 41.4 | 73.54 | |
| nl-ar | 3.1 | 18.2 | 67.72 | ru-en | 25.1 | 43.9 | 74.05 | |
| nl-de | 10.4 | 27.7 | 52.67 | zh-en | 32.6 | 54.3 | 79.75 | |
| nl-fr | 16.9 | 37.3 | 68.46 | avg. | 28.35 | 48.00 | 77.30 | |
| nl-ru | 4.8 | 17.8 | 54.71 | |||||
| nl-zh | 8.1 | 7.0 | 63.96 | |||||
| ru-ar | 11.9 | 31.5 | 72.45 | |||||
| ru-de | 6.1 | 23.7 | 52.74 | |||||
| ru-fr | 21.2 | 42.5 | 71.71 | |||||
| ru-nl | 6.8 | 22.6 | 53.91 | |||||
| ru-zh | 21.3 | 20.7 | 74.63 | |||||
| zh-ar | 13.1 | 34.1 | 74.92 | |||||
| zh-de | 4.1 | 22.3 | 52.13 | |||||
| zh-fr | 23.8 | 46.5 | 76.54 | |||||
| zh-nl | 4.8 | 19.9 | 54.26 | |||||
| zh-ru | 5.7 | 21.9 | 50.60 | |||||
| avg. | 11.79 | 26.36 | 63.22 | |||||
| seenseen | 22.68 | 35.32 | 76.39 | enseen | 26.20 | 37.77 | 78.22 | |
| seenunseen | 7.10 | 24.50 | 55.18 | enunseen | 16.17 | 33.27 | 63.50 | |
| unseenseen | 12.56 | 24.74 | 68.83 | seenen | 31.97 | 52.93 | 79.72 | |
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 6.7 | 24.2 | 57.45 | en-ar | 15.1 | 35.8 | 76.76 | |
| ar-fr | 27.5 | 49.2 | 75.21 | en-de | 20.6 | 35.9 | 65.88 | |
| ar-nl | 8.7 | 24.8 | 59.14 | en-fr | 27.5 | 49.4 | 77.46 | |
| ar-ru | 6.7 | 21.6 | 55.04 | en-nl | 15.0 | 35.6 | 63.70 | |
| ar-zh | 30.1 | 24.4 | 78.54 | en-ru | 13.5 | 29.5 | 61.62 | |
| de-ar | 3.5 | 19.7 | 68.39 | en-zh | 36.3 | 27.7 | 80.52 | |
| de-fr | 15.4 | 35.8 | 67.81 | avg. | 21.33 | 35.65 | 70.99 | |
| de-nl | 9.6 | 27.3 | 53.74 | |||||
| de-ru | 4.7 | 17.9 | 54.23 | |||||
| de-zh | 12.0 | 9.9 | 65.40 | |||||
| fr-ar | 14.9 | 36.3 | 74.98 | BLEU | chrF++ | COMET | ||
| fr-de | 9.2 | 28.3 | 52.96 | ar-en | 33.9 | 53.7 | 79.74 | |
| fr-nl | 11.3 | 31.1 | 57.62 | de-en | 27.1 | 43.6 | 77.13 | |
| fr-ru | 8.8 | 26.2 | 56.31 | fr-en | 29.7 | 51.0 | 80.03 | |
| fr-zh | 31.1 | 26.9 | 79.93 | nl-en | 22.6 | 42.3 | 73.94 | |
| nl-ar | 3.3 | 18.5 | 68.02 | ru-en | 25.8 | 44.5 | 74.07 | |
| nl-de | 9.4 | 26.5 | 52.33 | zh-en | 32.5 | 54.5 | 80.01 | |
| nl-fr | 17.2 | 37.3 | 68.38 | avg. | 28.60 | 48.27 | 77.49 | |
| nl-ru | 4.4 | 17.1 | 53.63 | |||||
| nl-zh | 8.3 | 7.0 | 64.08 | |||||
| ru-ar | 12.4 | 32.1 | 72.40 | |||||
| ru-de | 5.7 | 22.9 | 51.90 | |||||
| ru-fr | 21.5 | 42.7 | 72.08 | |||||
| ru-nl | 6.3 | 22.1 | 53.32 | |||||
| ru-zh | 22.7 | 24.6 | 75.36 | |||||
| zh-ar | 13.9 | 35.4 | 75.68 | |||||
| zh-de | 3.6 | 21.3 | 51.32 | |||||
| zh-fr | 24.5 | 47.0 | 76.98 | |||||
| zh-nl | 4.9 | 20.3 | 54.30 | |||||
| zh-ru | 5.5 | 21.1 | 50.49 | |||||
| avg. | 12.13 | 26.65 | 63.23 | |||||
| seenseen | 23.67 | 36.53 | 76.89 | enseen | 26.30 | 37.63 | 78.25 | |
| seenunseen | 7.27 | 24.32 | 54.96 | enunseen | 16.37 | 33.67 | 63.73 | |
| unseenseen | 12.92 | 25.29 | 69.10 | seenen | 32.03 | 53.07 | 79.93 | |